What is the capacitance of a parallel plate capacitor with metal plates, each of area of 6.00 m2, separated by 10.00 mm?


1
Expert's answer
2022-02-23T08:30:34-0500

The capacitance of a parallel plate capacitor is given by

"C=\frac{\epsilon_0 A}{d}\\
=\rm\frac{8.85*10^{-12}\: F/m*6.00\: m^2}{10.00*10^{-3}\: m}=5.31*10^{-9}\: F=5.31\: nF"